OpenStudy (anonymous):

mth help;// Brianna has 11 2/3 cups of milk. She needs 1 1/6 cups to makes a pot of hot cocoa. How many pots of hot cocoa can Brianna make? Write answer in simplest and sentence form. Mya played in two softball games one afternoon. The first game lasted 42 minutes. The second game lasted 1 2/3 times longer than the first game. How long did Mya's second game last? Write your answer in simplest and sentence form.

OpenStudy (danjs):

For the first one, you want to see how many times 1 1/6 can be divided into 11 2/3 can you change those to improper fractions first?

OpenStudy (anonymous):

ya

OpenStudy (danjs):

ok, do that first, then recall \[\large\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b }*\frac{ d }{ c }\]

OpenStudy (danjs):

what you think?>

OpenStudy (anonymous):

siiigh ill just do it myself nd stop being so lazy